# NoMIRACL: Knowing When You Don’t Know for Robust Multilingual Retrieval-Augmented Generation

### v2

@@ 5.1 NoMIRACL Non-relevant Subset
Figure 5 shows hallucination rates on the NoMIRACL non-relevant subset for a maximum of 250 queries evaluated (each language) on all 18 languages for best LLM in each category (for all model results, please refer to Figure 8). Our findings indicate that all LLMs (except Mistral) hallucinate that an answer is present across all languages, thereby indicating their poor ability to abstain from answering. On average, the lowest hallucination rate of 11.7% is observed by Mistral-7B, followed by Mixtral-7x8B with 17.4%. GPT-4 achieves a 33.2% hallucination rate, which highlights the challenge of identifying non-relevant passages. FLAN-T5, LLAMA-2, Orca-2 and Aya-101 perform much worse on average across all languages, by achieving a hallucination rate of more than 80%. We hypothesize that LLMs perform poorly to identify non-relevant passages as they are highly similar to the query, but do not contain the exact answer.
Overall, the lowest hallucination rates are observed in Swahili and Yoruba. We hypothesize that queries in low-resource languages (smaller Wikipedia corpus) contain retrieved information that is likely to be non-relevant (easier negative), thereby making it easier for the LLM to judge as “I don’t know”. GPT-3.5 (cf. Figure 8) is observed with the highest deviation across languages with a hallucination rate as low as 25.2% on Swahili (sw) to 95.2% on Bengali (bn). Overall, a majority of LLMs are found to perform poorly on NoMIRACL non-relevant subset, indicating our dataset is very challenging in robustness evaluation for LLMs.
@@ 5.2 NoMIRACL Relevant Subset
Figure 6 shows error rates on NoMIRACL relevant subset for a maximum of 250 queries (each language) on 18 languages for best LLM in each category (for all model results, please refer to Figure 9). Our findings indicate that all LLMs (except Mistral) are able to identify the answer present within the relevant passage accurately. On average, Mistral-7B achieves the highest error rate of 74.9%. The lowest error rates are observed by FLAN-T5, LLAMA-2-70B, and GPT-4 which are lower than 16%. Overall, Aya-101, Mixtral-7x8B and LLAMA-2-7B perform worse on average by achieving more than a 40% error rate. Overall, a all LMs (except Mistral and Aya-101) perform well and do not suffer from errors in identifying answers in the relevant subset.
@@ 5.3 NoMIRACL Overall Comparison
A robust LLM should be able to identify the answer captured within retrieved passages in the relevant subset and abstain from answering when none of the retrieved passages contain the answer in the non-relevant subset. To measure performance across both dimensions in NoMIRACL, we plot the average model accuracy across both the non-relevant (x-axis) and relevant subset (y-axis) for all tested models in Figure 7. Overall, LLMs positioned in the top-right corner provide an optimal performance on both subsets. A majority of LLMs (such as LLAMA-2, Orca-2, FLAN-T5) in the top-left corner perform well on the relevant subset, however, hallucinate and struggle to perform well on the non-relevant subset, indicating their inability to accurately determine non-relevant passages.
On the other hand, Mistral LLMs do not suffer from hallucination on the non-relevant subset but instead overall observe a higher error rate (over 40%) on the relevant subset, indicating they are not able to confidently identify passages containing the relevant answer. Aya-101 is unable to perform well in either of the subsets. GPT-4 overall provides the performance tradeoff balancing a low hallucination and error rate on NoMIRACL relevant and non-relevant subsets, however can be costly to be computed at scale for inference.
@@ 6 Empirical Analysis
In this section, we conduct an empirical analysis of LLM outputs, with respect to both the samples in the non-relevant subset (containing hallucinations) and relevant subset (containing errors) where LLMs fail to recognize answers within the relevant passages. We conduct our ablation on the English (en) subset in NoMIRACL. We categorized each LLM output pattern as either positive or accurate (highlighted in green), unable to understand instruction (highlighted in orange) or either a hallucination or error (highlighted in red).
Non-relevant subset. As shown in section 6 (above), we observe an uniform distribution of the hallucination pattern of failed samples for a majority of LLMs by answering “Yes, answer is present” with or without additional explanation. LLMs such as LLAMA-2-7B and LLAMA-2-13B suffer from hallucinations by often repeating the question or instruction in their generation output. Mistral-7B interestingly always provides a rationale or explanation in their model response, whereas Aya-101, uses implicit memory heavily to directly provide an answer instead of grounding the answer from within the retrieved passages. Lastly, models such as Orca-2-7B tend to change the output generation style and often use synonyms such as “No, answer is not present” instead of “I don’t know”.
Relevant subset. As shown in section 6 (below), similar to the non-relevant subset, we observe a uniform distribution of accurate LLM responses. Interestingly, GPT-4, Orca-2-13B and FLAN-T5-XXL overall only provide a single output classification token, whereas models such as GPT-3.5 or Mixtral-8x7B provide an additional rationale or explanation. Similar to the non-relevant subset, LLAMA-2-7B and 13B models repeat the question and instruction in their output and Aya-101 overall uses implicit memory to answer queries.
OpenAI Mistral Orca-2 Aya LLAMA-2 FLAN-T5
GPT-4 GPT-3.5 8x7B 7B 13B 7B Aya-101 70B 13B 7B XXL
@@
(i) Perfectly answers “I don’t know” 56.8% 54.8% 67.2% 1.6% 3.2% 9.2% 15.6% 0.0% 0.0% 0.8% 15.2%
(ii) “I don’t know” with explanation 0.0% 0.0% 3.6% 88.4% 0.0% 2.0% 0.0% 0.0% 0.0% 0.0% 0.0%
(iii) Uses a synonym of “I don’t know” 0.4% 0.0% 1.6% 1.2% 6.0% 7.6% 0.0% 0.4% 0.0% 2.4% 6.4%
@@
(v) Repeats question or instruction 0.0% 0.0% 0.0% 0.0% 0.0% 3.2% 0.0% 2.0% 64.8% 30.8% 0.0%
(vi) Conversation 0.0% 0.0% 0.4% 0.0% 0.0% 0.4% 0.0% 0.0% 0.0% 0.4% 0.0%
@@
(viii) Uses implicit memory to answer 0.0% 0.0% 0.0% 0.0% 0.0% 8.8% 10.8% 0.0% 0.0% 0.0% 0.4%
@@
(i) Perfectly answers “Yes, answer is present” 94.8% 45.2% 22.8% 5.2% 97.2% 71.2% 51.6% 1.2% 3.6% 7.2% 97.6%
(ii) "Yes, answer is present" with explanation 0.0% 46.0% 56.0% 31.6% 0.8% 11.2% 0.0% 98.8% 46.8% 63.2% 0.0%
@@
(iv) Repeats question or instruction 0.0% 0.0% 0.4% 0.0% 0.0% 0.4% 0.0% 0.0% 48.8% 24.8% 0.0%
(v) Conversation 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.6% 0.0%
@@
(vii) Uses implicit memory to answer 0.0% 0.0% 0.0% 2.4% 1.2% 14.4% 42.0% 0.0% 0.0% 0.4% 0.0%
@@ 7 Related Work
Retrieval-Augmented Generation. The knowledge stored in a large language model (LLM) is commonly outdated He et al. (2023), and prone to hallucinations by generating factually incorrect output Maynez et al. (2020); Raunak et al. (2021). By grounding on external knowledge, a retrieval-augmented LLM can generate better and more trustworthy output Guu et al. (2020); Lewis et al. (2020); Izacard and Grave (2021); Borgeaud et al. (2022). Retrieval-augmented generation has achieved remarkable results in various tasks such as open-domain question answering (ODQA) Lewis et al. (2020); Izacard and Grave (2021); Trivedi et al. (2023), argument extraction Du and Ji (2022) and code generation Zhou et al. (2023). Real-world products such as Bing Search [^6] New Bing Search: news.microsoft.com/the-new-Bing/ and Langchain [^7] Langchain: github.com/langchain-ai/langchain have incorporated RAG applications.
@@
Related Datasets. Datasets focused on addressing unanswerable queries such as SQuAD 2.0 Rajpurkar et al. (2018) were created in an adversarial way to look similar to datasets with answerable queries. Similarly, Conversational QA datasets such as CoQA Reddy et al. (2019) and QuAC Choi et al. (2018) also contain unanswerable queries. A concurrent work proposes RGB, a RAG benchmark to evaluate LLM robustness in English and Chinese Chen et al. (2023).
@@ 8 Conclusion
We introduce NoMIRACL, a human-annotated hallucination dataset for evaluating the performance of LLM robustness in recognizing no-answers as a binary classification task in 18 languages. Our multilingual dataset is human-annotated and constructed with 31 native speakers. We provide two subsets in NoMIRACL, the non-relevant subset, where queries contain all judged non-relevant passages, and the relevant subset, where queries contain at least a single relevant judged passage to measure both hallucination on the non-relevant and error on the relevant subset. Our experimental results indicate that existing LLMs are not robust, as we observe challenges in LLM robustness in either hallucination or error. Overall, GPT-4 achieves the best model and performance tradeoff across both subsets. We observed several hallucination patterns within LLMs. NoMIRACL can facilitate research in understanding to which extent LLMs tend to hallucinate, ultimately paving the way for building more effective and robust LLMs in the future.
@@ 9 Limitations
NoMIRACL is not perfect and like other datasets have limitations. We describe our limitations below and keep it as future work to continue to improve our dataset.
1. No finetuning. In our work, we focus on robustness evaluation on a wide variety of multilingual LLMs. We keep as future work to explore supervised fine-tuning LLMs to reduce hallucinations in LLMs, such as Chain-of-Verification (CoVe; Dhuliawala et al., 2023) or recently introduced Chain-of-Noting (CoN; Yu et al., 2023).
2. Human errors in dataset construction. Our dataset has been fully constructed using humans, thereby it may contain human errors. We conducted an additional quality checks on a subset of the NoMIRACL dataset to validate its question quality and relevance judgement as explained in Appendix C.
3. Evaluation setup. In our work, we evaluate whether a passage is either relevant or non-relevant passage for a given query, instead of evaluating actual answer spans. Coming up with reliable and accurate answers for a given query requires domain experts as annotators. Annotators can potentially highlight short extractive spans of answers within relevant passages, however, non-extractive queries can either contain multiple answers or a long-form answer, making it difficult to highlight a relevant answer span. Therefore, for NoMIRACL, we focus on the evaluation of top-$k$ passages as information contexts, which were easily judged for their relevancy by a data annotator.
4. Prompt Optimization. In our experiments, we evaluate models using a single vanilla prompt template. We leave it as future work, to further experiment with better and prompt optimzation techniques, such as including Chain-of-Thought (CoT; Wei et al., 2022).

### v3

@@ 5.1 NoMIRACL Non-relevant Subset
Figure 5 shows hallucination rates on the NoMIRACL non-relevant subset for a maximum of 250 queries evaluated (each language) on all 18 languages for best LLM in each category (for all model results, please refer to Figure 8). Our findings indicate that all LLMs (except Mistral) hallucinate that an answer is present across all languages, thereby indicating their poor ability to abstain from answering. On average, the lowest hallucination rate of 17.4% is observed by Mixtral-7x8B, followed by LLAMA-3-8B-Instruct with 26.8%. GPT-4 achieves a 35.5% hallucination rate, which highlights the challenge of identifying non-relevant passages. LLAMA-2, Orca-2, and Aya-101 perform much worse on average across all languages, by achieving a hallucination rate of more than 80%. We hypothesize that LLMs perform poorly to identify non-relevant passages as they are highly similar to the query, but do not contain the exact answer.
Overall, the lowest hallucination rates are observed in Swahili and Yoruba. We hypothesize that queries in low-resource languages (smaller Wikipedia corpus) contain retrieved information, likely to be non-relevant (easier negative), thereby making it easier for the LLM to judge as “I don’t know”. GPT-3.5 (cf. Figure 8) is observed with the highest deviation across languages with a hallucination rate as low as 25.2% on Swahili (sw) to 95.2% on Bengali (bn). Overall, all LLMs are found to perform poorly on NoMIRACL non-relevant subset, indicating our dataset is very challenging in robustness evaluation for LLMs.
@@ 5.2 NoMIRACL Relevant Subset
Figure 6 shows error rates on NoMIRACL relevant subset for a maximum of 250 queries (each language) on 18 languages for best LLM in each category Our findings indicate that all LLMs (except Mistral) identify the answer present within the relevant passage. On average, Aya-101 achieves the highest error rate of 62.5%. The lowest error rates are observed by LLAMA-2-70B and GPT-4o which are lower than 10%. Overall, Aya-101, Mixtral-7x8B, and LLAMA-3-8B perform worse on average by observing more than a 40% error rate. Overall, LLMs (except Mistral and Aya-101) perform well and do not suffer from errors in identifying answers in the NoMIRACL relevant subset.
@@ 5.3 NoMIRACL Overall Comparison
A robust LLM should be able to identify the answer captured within retrieved passages in the relevant subset and abstain from answering when none of the retrieved passages contain the answer in the non-relevant subset. To measure performance across both dimensions in NoMIRACL, we plot the average model accuracy across both the non-relevant (x-axis) and relevant subset (y-axis) for all tested models in Figure 7. Overall, LLMs positioned in the top-right corner provide an optimal performance on both subsets. A majority of LLMs (such as LLAMA-2, and Orca-2) in the top-left corner perform well on the relevant subset, however, hallucinate and struggle to perform well on the non-relevant subset, indicating their inability to accurately judge non-relevant passages.
On the other hand, Mistral and LLAMA-3 suffer less from hallucination on the non-relevant subset but observe a higher error rate (over 40%) on the relevant subset, indicating they are not confident in identifying passages containing the answer. Aya-101 is unable to perform well in either of the subsets. GPT-4 provides a good tradeoff balancing both a low hallucination and error rate on NoMIRACL relevant and non-relevant subsets, however is expensive to compute at scale for inference.
@@ 6 Empirical Analysis
In this section, we conduct an empirical analysis of LLM outputs, with both the non-relevant subset (containing hallucinations) and the relevant subset (containing errors). We conduct our ablation study on the English (en) subset in NoMIRACL. We categorize each LLM output pattern as either positive or accurate (highlighted in green), unable to understand instruction (highlighted in orange), or either a hallucination or error (highlighted in red).
Non-relevant Subset. As shown in section 6 (above), we observe a uniform distribution of the hallucination pattern of failed samples for a majority of LLMs by answering “Yes, answer is present” with or without additional explanation. LLMs such as LLAMA-2-7B and LLAMA-2-13B suffer from hallucinations by often repeating the question or instruction in their generation output. Mistral-7B interestingly always provides a rationale or explanation in their model response, whereas Aya-101, uses implicit memory heavily to directly provide an answer instead of grounding the answer from within the retrieved passages. Lastly, models such as Orca-2-7B tend to change the output generation style and often use synonyms such as “No, answer is not present” instead of “I don’t know”.
Relevant Subset. As shown in section 6 (below), similar to the non-relevant subset, we observe a uniform distribution of accurate LLM responses. Interestingly, GPT-4 and Orca-2-13B overall only provide a single output classification token, whereas models such as GPT-3.5 or Mixtral-8x7B provide an additional rationale or explanation. Similar to the non-relevant subset, LLAMA-2-7B and 13B models repeat the instruction in their output and Aya-101 sometimes uses implicit memory.
OpenAI Mistral Orca-2 Aya LLAMA-2
GPT-4 GPT-3.5 8x7B 7B 13B 7B Aya-101 70B 13B 7B
@@
(i) Perfectly answers “I don’t know” 56.8% 54.8% 67.2% 1.6% 3.2% 9.2% 15.6% 0.0% 0.0% 0.8%
(ii) “I don’t know” with explanation 0.0% 0.0% 3.6% 88.4% 0.0% 2.0% 0.0% 0.0% 0.0% 0.0%
(iii) Uses a synonym of “I don’t know” 0.4% 0.0% 1.6% 1.2% 6.0% 7.6% 0.0% 0.4% 0.0% 2.4%
@@
(v) Repeats question or instruction 0.0% 0.0% 0.0% 0.0% 0.0% 3.2% 0.0% 2.0% 64.8% 30.8%
(vi) Conversation 0.0% 0.0% 0.4% 0.0% 0.0% 0.4% 0.0% 0.0% 0.0% 0.4%
@@
(viii) Uses implicit memory to answer 0.0% 0.0% 0.0% 0.0% 0.0% 8.8% 10.8% 0.0% 0.0% 0.0%
@@
(i) Perfectly answers “Yes, answer is present” 94.8% 45.2% 22.8% 5.2% 97.2% 71.2% 51.6% 1.2% 3.6% 7.2%
(ii) "Yes, answer is present" with explanation 0.0% 46.0% 56.0% 31.6% 0.8% 11.2% 0.0% 98.8% 46.8% 63.2%
@@
(iv) Repeats question or instruction 0.0% 0.0% 0.4% 0.0% 0.0% 0.4% 0.0% 0.0% 48.8% 24.8%
(v) Conversation 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 1.6%
@@
(vii) Uses implicit memory to answer 0.0% 0.0% 0.0% 2.4% 1.2% 14.4% 42.0% 0.0% 0.0% 0.4%
@@
7 Further Studies
Prompt Optimization. Prompting is crucial in handling the robustness evaluation of multilingual-focused LLMs. Techniques such as Chain-of-Thought (CoT) Wei et al. (2022) or algorithmically optimizing prompts using DSPy Khattab et al. (2023) highlight the necessity of prompt optimization. Although optimizing for the prompt is certainly challenging and expensive to evaluate all LLMs across 18 languages relevant and non-relevant subsets, we experiment with three listwise variations techniques inspired by Thomas et al. (2024). The prompt template changes are listed in section 6: (i) role, we highlight the role of LLM as an evaluator within the prompt at the beginning, (ii) repeat, we repeat the task instructions at the end of the prompt to remind the LLM, and (iii) explanation, we ask the LLM model to provide a step-by-step explanation and then answer and require 400 output tokens to fit both the LLM reasoning and the answer.
We evaluate Mistral-7B with three prompt variations independently on NoMIRACL. The complete results are listed in Table 3. On average, both role and repeat techniques help reduce the error rate in the NoMIRACL relevant subset by 6.3% and 15.2% but overall increase the hallucination rate by 8.7% and 15.9% respectively. On the other hand, prompting with explanation decreases the hallucination rate by 9.7% but increases the error rate by 8.3%. These results show that prompting is user dependent, the user will be required to choose their technique depending on whether they wish to be better on the non-relevant subset by reducing the hallucination rate or the relevant subset by reducing the error rate.
ar bn de en es fa fr fi hi id ja ko ru sw te th yo zh Avg.
Hallucination Rates (in %) on NoMIRACL test split (non-relevant subset)
Original 40.0 63.2 38.2 42.8 17.2 52.4 47.6 16.1 39.6 30.8 44.4 28.8 41.6 14.8 74.0 58.8 23.6 45.2 40.0
(+ Role) 35.6 60.0 53.5 62.4 40.4 38.0 71.2 29.8 38.4 49.2 59.2 41.2 55.6 21.2 72.0 54.4 32.4 62.8 48.7
(+ Repeat) 47.2 72.4 56.7 50.8 35.2 69.2 70.0 50.0 49.6 48.8 65.2 48.4 56.0 35.2 78.4 72.0 43.6 58.4 55.9
(+ Explanation) 28.0 33.6 30.4 34.8 16.8 39.6 42.4 31.5 18.0 27.2 32.8 26.8 31.6 22.8 36.4 34.0 27.2 31.6 30.3
Error Rates (in %) on NoMIRACL test split (relevant subset)
Original 14.4 20.4 21.6 8.0 28.0 22.4 15.6 38.0 30.8 43.6 16.0 30.0 18.0 57.6 17.6 13.6 50.0 13.2 25.5
(+ Role) 15.2 23.6 10.0 3.2 14.4 23.2 7.2 29.2 25.2 32.8 9.6 21.6 10.8 47.2 17.6 17.2 33.8 4.4 19.2
(+ Repeat) 8.4 12.4 9.2 5.2 12.4 4.8 5.6 6.0 16.4 16.8 1.6 12.0 6.8 22.8 14.0 6.0 21.6 3.6 10.3
(+ Explanation) 28.8 39.6 26.0 16.8 35.6 35.6 20.0 32.0 46.8 46.8 24.4 33.6 24.8 50.4 48.8 31.2 48.0 18.8 33.8
Table 3: Hallucination and error rates on the NoMIRACL test split (non-relevant and relevant subsets) with three types of prompting techniques on Mistral-7B (v0.3). The changes in the prompt template are listed in section 6.
Fine-tuning on NoMIRACL. In this section, following prior works such as Chain-of-Verification (CoVe; Dhuliawala et al., 2024) or Chain-of-Noting (CoN; Yu et al., 2023), we investigate the following research question: Does fine-tuning on the NoMIRACL development set help increase robustness?
We experiment with two open-sourced LLMs: Mistral-7B and LLAMA-3 (8B). We Supervised Fine-Tune (SFT) LoRA adapters Hu et al. (2022) on the development set of NoMIRACL for all 18 languages (randomly sampled 90% train, 10% development) using 4-A6000 GPUs each containing 48GB RAM with PEFT. [^5] https://github.com/huggingface/alignment-handbook Our hyperparameter settings are listed in Table 6. We were unable to fine-tune larger models (greater than 8B parameters) due to computational budget restrictions.
As shown in Table 4, we observe LLAMA-3 (8B) to be quite unstable after SFT. Fine-tuning helps to reduce the error rate of LLAMA-3 (8B) (an improvement of 10.6%) but can hurt its performance on the hallucination rate (drop up to 17.9%). For a few languages mentioned in Table 7 such as Arabic (ar) the LLM always outputs “Yes, answer is present”, whereas for Bengali (bn) heavily relies on “I don’t know”. On the other hand, SFT deteriorates Mistral-7B on both relevant and non-relevant datasets. Overall, we demonstrate SFT is tricky and careful experimentation is required to achieve the best out of fine-tuning on the NoMIRACL development subset for a binary classification task output (“Yes, answer is present” or “I don’t know”).
Model w/o SFT w/ SFT
Non-Relevant Subset: Hallucination Rates (in %)
Meta-Llama-3-8B-Instruct 26.8 44.7 (– 17.9)
Mistral-Instruct-7B-v0.3 40.0 44.3 (– 4.3)
Relevant Subset: Error Rates (in %)
Meta-Llama-3-8B-Instruct 45.3 34.7 (+ 10.6)
Mistral-Instruct-7B-v0.3 25.5 46.1 (– 20.6)
Table 4: Supervised fine-tuning on the NoMIRACL development split with Llama-3 (8B) and Mistral-7B (v0.3) LLMs.
@@ 8 Related Work
Retrieval-Augmented Generation. The knowledge stored in a large language model (LLM) is commonly outdated He et al. (2023), and prone to hallucinations by generating factually incorrect output Maynez et al. (2020); Raunak et al. (2021). By grounding on external knowledge, a retrieval-augmented LLM can generate better and more trustworthy output Guu et al. (2020); Lewis et al. (2020); Izacard and Grave (2021); Borgeaud et al. (2022). Retrieval-augmented generation has achieved remarkable results in various tasks such as open-domain question answering (ODQA) Lewis et al. (2020); Izacard and Grave (2021); Trivedi et al. (2023), argument extraction Du and Ji (2022) and code generation Zhou et al. (2023). Real-world products such as Bing Search and LangChain have incorporated RAG applications.
@@
Related Datasets. Datasets focused on addressing unanswerable queries such as SQuAD 2.0 Rajpurkar et al. (2018) were created adversarially to look similar to datasets with answerable queries. Similarly, Conversational QA datasets such as CoQA Reddy et al. (2019) and QuAC Choi et al. (2018) also contain unanswerable queries. A concurrent work proposes RGB, a RAG benchmark to evaluate LLM robustness in English and Chinese Chen et al. (2024).
@@ 9 Conclusion
We introduce NoMIRACL, a multilingual human-labeled dataset for relevance assessment of LLM robustness as a binary relevance identification task in 18 languages. Our multilingual dataset is human-annotated and constructed with 31 native speakers. We provide two subsets in NoMIRACL, the non-relevant subset, where queries contain all judged non-relevant passages, and the relevant subset, where queries contain at least one relevant judged passage to measure the hallucination on the non-relevant and error on the relevant subset. Our experimental results indicate that existing LLMs are not robust, as we observe challenges in LLM robustness in either hallucination or error. GPT-4 achieves the best model and performance tradeoff across both subsets. NoMIRACL can facilitate research in understanding to which extent LLMs tend to hallucinate, ultimately paving the way for building more effective and robust multilingual-focused LLMs in the future.
@@ 10 Limitations
NoMIRACL is not perfect and like other datasets have limitations. We describe our limitations below and keep it as future work to improve our dataset.
1. Human Errors in Dataset Construction. Our dataset has been fully constructed using humans, thereby it may contain human errors. We conducted additional quality checks on a subset of the NoMIRACL dataset to validate its question quality and relevance judgment as explained in Appendix C.
2. Evaluation Setup. In our work, we evaluate whether a passage is relevant or non-relevant for a given query, instead of evaluating actual answer spans. Reliable and accurate answers for a given query require domain experts as annotators. Annotators can potentially highlight short extractive spans of answers within relevant passages, however, non-extractive queries can either contain multiple answers or a long-form answer, making it difficult to highlight a relevant answer span. Therefore, for NoMIRACL, we focus on evaluating top-$k$ passages as information contexts, which are judged for their relevancy by a data annotator.
3. Limited to Wikipedia. NoMIRACL is currently developed using language-specific Wikipedia as the corpora. Wikipedia may not be the ideal choice for real-world applications across languages. For example, the English BEIR benchmark Thakur et al. (2021) includes diversity within its domains (all English) and contains more real-world domains such as Medical, etc. However, we keep it as future work to extend NoMIRACL to diverse domains for the following reasons: (i) scarcity of corpora across languages: for low-resource languages such as Bengali or Yoruba, finding a suitable large enough text corpora is difficult with limited choices. (ii) no uniformity across domains: certain European languages have more legal domain corpora available, whereas news articles for African languages. This will introduce non-uniformity in information across languages. (iii) limited budget: constructing NoMIRACL was expensive involving several annotators involved for about 4–6 months. Extending to more domains would require additional budgets and human effort to be able to implement.